# Talha Bozkus and Urbashi Mitra

### v2

@@ VI-C Numerical consistency of propositions
Fig.5d illustrates the simulation misdetection probability ($P_{mis}$) vs. ARSS threshold $I_{thr}$, which varies in [$I_{min},I_{max}$], along with the upper and lower bounds from Proposition 4 for $N_{T}=2$, and the coordination percentage (i.e. $\frac{|\mathcal{S}_{c}|}{|\mathcal{S}|}$), where the blue shaded area represents the standard deviation over 200 simulations. We make several observations. First, the upper and lower bounds in Proposition 4 hold for any ARSS threshold value. Second, there is a unique ARSS threshold that minimizes the $P_{mis}$, which can also be theoretically obtained in Proposition 4. Third, as the ARSS threshold increases, the proportion of coordinated states decreases monotonically. When $P_{mis}$ is minimized, coordinated states make up around 35%, indicating that even in the optimal scenario, the system largely operates independently. Fourth, $P_{mis}$ is larger when $I_{thr}$ is very small (close to the minimum ARSS measurement) than when it is very large (close to the maximum ARSS measurement). This is expected as large $I_{thr}$ causes most states to be uncoordinated, reducing the likelihood of misdetection. Similarly, even though $I_{thr}$ changes significantly, the change in $P_{mis}$ remains relatively modest, demonstrating the robustness of the system to change in system parameters as in Table IV.
@@ VII Conclusions
In this paper, we present a novel multi-agent MEMQ algorithm for cooperative decentralized wireless networks with multiple networked transmitters (TXs) and base stations (BSs) to address the performance and complexity limitations of single-agent MEMQ and existing MARL algorithms. Transmitters lack access to global information but can locally measure ARSS due to other TXs. TXs act independently to minimize their individual costs in uncoordinated states. In coordinated states, they employ a Bayesian approach to estimate the joint state based on local ARSS measurements and share limited information with only the leader TX to minimize the joint cost. This communication scheme incurs an information-sharing cost that scales only linearly with the number of TXs, independent of the joint state-action space size. We provide several deterministic and probabilistic theoretical results, including convergence, upper bounds on estimation error variance, and state misdetection probability. We numerically show that our algorithm outperforms several decentralized and CTDE algorithms by achieving 55% lower APE with 50% lower runtime complexity and 35% faster convergence with 45% less sample complexity. It significantly reduces the complexity of centralized algorithms while achieving comparable APE. In the end, it is shown that the simulation results closely follow the theoretical results. For our future work, we work on developing a theoretical sample complexity expression to enable direct comparison with MARL algorithms. We also want to address the continuous state-spaces by combining appropriate deep RL algorithms with our algorithm. Finally, we focus on devising computationally efficient sampling methods to ensure good data coverage across all environments and all agents as in [36, 37].
-A Proof of Proposition 2
We first rewrite (4) as follows:
$Q^{e}_{i,t}=uQ^{e}_{i,t}+(1-u)\sum_{n=1}^{K_{i}}w_{i,t}^{(n)}Q_{i,t}^{(n)},$ (24)
where we drop the $(s_{i},a_{i})$ notation for simplicity and introduce time subscript $t$. By repeatedly plugging the expression of $Q^{e}_{i,t-1}$ into the expression of $Q^{e}_{i,t}$ in (24), the following can be shown:
$Q^{e}_{i,t}$ $=(1-u)\sum_{i=0}^{t-1}u^{t-i-1}\sum_{n=1}^{K}w_{i,t}^{(n)}Q_{i,t}^{(n)}.$ (25)
Using (25) and algebraic manipulations, we can show:
$Q^{e}_{i,t}-uQ^{e}_{i,t-1}=(1-u)\sum_{n=1}^{K}w_{i,t-1}^{(n)}Q_{i,t-1}^{(n)}.$ (26)
Let $\Delta_{i,t}=Q_{i,t+1}^{e}-Q_{i,t}^{e}$, $\epsilon_{i,t}^{(n)}=w_{i,t+1}^{(n)}Q_{i,t+1}^{(n)}-w_{i,t}^{(n)}Q_{i,t}^{(n)}$. If we rewrite (26) with the variable change $t\rightarrow t+1$, and subtract (26) from the new expression side by side, we obtain:
$\Delta_{i,t}-u\Delta_{i,t-1}^{it}$ $=(1-u)\sum_{n=1}^{K}\Big{[}w_{i,t}^{(n)}Q_{i,t}^{(n)}-w_{i,t-1}^{(n)}Q_{i,t-1}^{(n)}\Big{]}.$ (27)
Then, we can show the following:
$\Delta_{i,t}$ $=u\Delta_{i,t-1}+(1-u)\sum_{n=1}^{K}\epsilon_{i,t-1}^{(n)}.$ (28)
$<\Delta_{i,t-1}+(1-u)\sum_{n=1}^{K}\epsilon_{i,t-1}^{(n)},$ (29)
where (28) follows from (27) and the definition of $\epsilon_{i,t-1}^{(n)}$ and (29) follows from the fact that $u\in(0,1)$. Then, using (29) and the triangle inequality, we can bound the difference between consecutive updates as follows:
$|\Delta_{i,t}-\Delta_{i,t-1}|$ $<(1-u)\sum_{n=1}^{K}|\epsilon_{i,t-1}^{(n)}|.$ (30)
Herein, $\lim_{t\rightarrow\infty}Q_{i,t-1}^{(n)}=Q_{i,t}^{(n)}$ by the convergence of $Q$-learning [35] and $\lim_{t\rightarrow\infty}w_{i,t-1}^{(n)}=w_{i,t}^{(n)}$ by the convergence of the weights [6] for all $i,n$. Thus, $\lim_{t\rightarrow\infty}|\epsilon_{i,t-1}^{(n)}|=0$ for all $i,n$, and hence:
$\lim_{t\rightarrow\infty}|\Delta_{i,t}-\Delta_{i,t-1}|=0,\quad\forall i.$ (31)
Using the definition of $\bar{\Delta}_{t}(\bar{s},\bar{a})$, we can show:
$\bar{\Delta}_{t}(\bar{s},\bar{a})$ $=\bar{Q}_{t}(\bar{s},\bar{a})-\bar{Q}_{t-1}(\bar{s},\bar{a}).$ (32)
$=\sum_{i=1}^{N_{T}}(Q_{i,t}^{e}(s_{i},a_{i})-Q_{i,t-1}^{e}(s_{i},a_{i})).$ (33)
$=\sum_{i=1}^{N_{T}}\Delta_{i,t}(s_{i},a_{i}),$ (34)
where (33) follows from the update rule in Table II and (34) follows from the definition of $\Delta_{i,t}$. Then:
$\lim_{t\rightarrow\infty}|\bar{\Delta}_{t}$ $(\bar{s},\bar{a})-\bar{\Delta}_{t-1}(\bar{s},\bar{a})|=$
$=\lim_{t\rightarrow\infty}|\sum_{i=1}^{N_{T}}(\Delta_{i,t}(s_{i},a_{i})-\Delta_{i,t-1}(s_{i},a_{i}))|,$ (35)
$\leq\sum_{i=1}^{N_{T}}\lim_{t\rightarrow\infty}(\Delta_{i,t}(s_{i},a_{i})-\Delta_{i,t-1}(s_{i},a_{i})),$ (36)
$=0,$ (37)
where (35) follows from (34), (36) follows from the linearity of expectation and triangle inequality, (37) follows from (31), where we introduce the notation $(s_{i},a_{i})$ back.
-B Proof of Proposition 4
The probability of detection error is the sum of probabilities for misdetecting a coordinated state as uncoordinated and vice versa. Let $\bar{s}$ be the actual state, and $\hat{s}$ be the detected state.
$P_{\text{err}}$ $=P(\bar{s}=\text{U},\hat{s}=\text{C})+P(\bar{s}=\text{C},\hat{s}=\text{U}).$ (38)
$=P(\bar{s}=\text{U}\mid\hat{s}=\text{C})P(\hat{s}=\text{C})+$
$\hskip 90.0ptP(\bar{s}=\text{C}\mid\hat{s}=\text{U})P(\hat{s}=\text{U}).$ (39)
$=P(I_{t}<I_{thr}\mid\hat{s}=\text{C})\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+$
$\hskip 90.0ptP(I_{t}>I_{thr}\mid\hat{s}=\text{U})\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|}.$ (40)
$=Q\left(\frac{I_{t,\text{true}}-I_{\text{thr}}}{\sigma_{c}}\right)\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+Q\left(\frac{I_{\text{thr}}-I_{t,\text{true}}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|},$ (41)
where (38) follows from the definition of misdetection, (39) follows from the Bayes rule, (40) follows from the ARSS criteria used to determine the state, and (41) follows from the noisy ARSS model in (17). We take the gradient of (41) with respect to $I_{\text{thr}}$ to find the critical points:
$\frac{dP_{\text{err}}}{dI_{\text{thr}}}$ $=\frac{1}{\sigma_{c}}Q^{\prime}\left(\frac{I_{t,\text{true}}-I_{\text{thr}}}{\sigma_{c}}\right)\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}-$
$\hskip 90.0pt\frac{1}{\sigma_{u}}Q^{\prime}\left(\frac{I_{\text{thr}}-I_{t,\text{true}}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|}.$ (42)
$=-\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|\sigma_{c}\sqrt{2\pi}}e^{-\frac{(I_{t,\text{true}}-I_{\text{thr}})^{2}}{2\sigma_{c}^{2}}}+$
$\hskip 80.0pt\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|\sigma_{u}\sqrt{2\pi}}e^{-\frac{(I_{\text{thr}}-I_{t,\text{true}})^{2}}{2\sigma_{u}^{2}}}.$ (43)
Setting (43) to zero, and taking the natural logarithm on both sides, we obtain:
$I_{\text{thr}}$ $=I_{t,\text{true}}\pm\sqrt{2\ln\left(\frac{|\mathcal{S}_{C}|\sigma_{u}}{|\mathcal{S}_{U}|\sigma_{c}}\right)\left(\frac{1}{\sigma_{c}^{2}}-\frac{1}{\sigma_{u}^{2}}\right)^{-1}}.$ (44)
If we plug this expression into (41), we obtain:
$P_{\text{err}}\geq Q\left(\frac{\Delta_{I}}{\sigma_{c}}\right)\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+Q\left(\frac{-\Delta_{I}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|}.$ (45)
On the other hand, if $I_{t,\text{true}}-I_{\text{thr}}<0$, then $\frac{d^{2}P_{\text{err}}}{dI_{\text{thr}}^{2}}<0$, indicating that (41) is concave in $I_{\text{thr}}$ and $I_{\text{thr}}=I_{t,\text{true}}+\Delta_{I}$ is the global maximum, as $\Delta_{I}>0$. If we plug this into (41), we obtain:
$P_{\text{err}}\leq Q\left(\frac{-\Delta_{I}}{\sigma_{c}}\right)\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+Q\left(\frac{\Delta_{I}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|}.$ (46)

### v3

@@ VI-C Numerical consistency of propositions
Fig.5d illustrates the simulation misdetection probability ($P_{mis}$) vs. ARSS threshold $I_{thr}$, which varies in $I_{min},I_{max}$], along with the upper and lower bounds from Proposition 4 for $N_{T}=2$, and the coordination percentage (i.e. $\frac{|\mathcal{S}_{c}|}{|\mathcal{S}|}$), where the blue shaded area represents the standard deviation over 200 simulations. We make several observations. First, the upper and lower bounds in Proposition 4 hold for any ARSS threshold value. Second, there is a unique ARSS threshold that minimizes the $P_{mis}$, which can also be theoretically obtained in Proposition 4. Third, as the ARSS threshold increases, the proportion of coordinated states decreases monotonically. When $P_{mis}$ is minimized, coordinated states make up around 35%, indicating that even in the optimal scenario, the system largely operates independently. Fourth, $P_{mis}$ is larger when $I_{thr}$ is very small (close to the minimum ARSS measurement) than when it is very large (close to the maximum ARSS measurement). This is expected as large $I_{thr}$ causes most states to be uncoordinated, reducing the likelihood of misdetection. Similarly, even though $I_{thr}$ changes significantly, the change in $P_{mis}$ remains relatively modest, demonstrating the robustness of the system to change in system parameters as in Table IV.
@@ VII Conclusions
We present a novel multi-agent MEMQ algorithm for cooperative decentralized wireless networks with multiple networked transmitters (TXs) and base stations (BSs) to address the performance and complexity limitations of single-agent MEMQ and existing MARL algorithms. Transmitters lack access to global information but can locally measure ARSS due to other TXs. TXs act independently to minimize their individual costs in uncoordinated states. In coordinated states, they employ a Bayesian approach to estimate the joint state based on local ARSS measurements and share limited information with only the leader TX to minimize the joint cost. This communication scheme incurs an information-sharing cost that scales only linearly with the number of TXs, independent of the joint state-action space size. We provide several deterministic and probabilistic theoretical results, including convergence, upper bounds on estimation error variance, and state misdetection probability. We numerically show that our algorithm outperforms several decentralized and CTDE algorithms by achieving 60% lower APE with 45% lower runtime complexity and 40% faster convergence with 40% less sample complexity. It significantly reduces the complexity of centralized algorithms while achieving comparable APE. It is shown that the simulation results closely follow the theoretical results. For our future work, we work on developing a theoretical sample complexity expression to enable direct comparison with MARL algorithms. Our algorithm is not directly applicable to continuous state–action spaces. To this end, we are extending the tabular multi-agent MEMQ framework to deep RL using an actor–critic architecture. Finally, we focus on devising efficient sampling methods to ensure good data coverage across all environments and all agents as in [34, 35].